How far do you go in two minutes at 18 mph
Answer by stanbon(75887)

How far do you go in two minutes at 18 mph
-----
distance = rate*time
---
distance = (18miles/hour)(2 minutes)
-----
= (18 miles/hour)(2 minutes)(1 hr/60 minutes)
----
= (18*2*1/1*1*60)(miles)
----
= (36/60) miles
----
= 3/5 miles
================
Cheers,
Stan H.
=============